OpenStudy (anonymous):

how do i integrate sqrt(x^2-49)/x^4

OpenStudy (luigi0210):

@modphysnoob quotient rule right?

OpenStudy (anonymous):

I know I have to use trig substitution, but not sure how

OpenStudy (anonymous):

this is integration , quotient rule is for differentiation

OpenStudy (luigi0210):

oops, sorry I haven't done calc in over 2 months :/ But yea now I remember

OpenStudy (anonymous):

this probably involves something way more than trig sub

OpenStudy (anonymous):

i get stuck after i make the substitution using secant

OpenStudy (anonymous):

i know i have to be familiar with sohcahtoa

OpenStudy (luigi0210):

Can't you rewrite and work backwards from there? \[\int\limits (x^2-49)^{1/2}x^{-4}dx\]

OpenStudy (amistre64):

you have to use trig? if so, then id go with either a sin or cos setup

OpenStudy (amistre64):

or tangent/cotangnet stuff

OpenStudy (amistre64):

sec^2 - 1 = tan^2

OpenStudy (anonymous):

for this particular problem I have make a substitution using secant

OpenStudy (amistre64):

let x = 7sec(t)

OpenStudy (anonymous):

I have that but after i plug everything in I get a little lost

OpenStudy (anonymous):

i am stuck at sqrt (7sec(t))^2/(7 sec(t))^4 * 7sec(t)tan(t)

OpenStudy (amistre64):

x = 7sec(t) dx = 7sec(t)tan(t) dt \[\frac{\sqrt{(x)^2-49}}{(x)^4}~(dx)\] replace parts \[\frac{\sqrt{(7sec(t))^2-49}}{(7sec(t))^4}~(7sec(t)tan(t)~dt)\] \[\frac{\cancel{7sec(t)}tan(t)~\sqrt{49}\sqrt{sec^2(t)-1}}{7^3 sec^3(t)}~~dt\] \[\frac{7sec(t)~tan(t)~tan(t)}{7^3 sec^3(t)}~~dt\] \[\frac{tan^2(t)}{7^2 sec^2(t)}~~dt\]

OpenStudy (amistre64):

\[tan=\frac{sin}{cos}\] \[\frac1{sec}=cos\] \[\frac{tan}{sec}=\frac{sin}{cos}cos=sin\]

OpenStudy (amistre64):

\[\frac1{49}\int sin^2(t)~dt\]

OpenStudy (amistre64):

theres a pretty simple trig identity for sin^2: 1/2 - cos(2t)/2

OpenStudy (anonymous):

I see. So i wouldn't need to draw a triangle because my professor showed us examples using sohcahtoa

OpenStudy (amistre64):

the triangle would be useful for after the integration to compare it back to a function without trigs

OpenStudy (anonymous):

I understand. So the triangle isn't absolutely necessary ?

OpenStudy (amistre64):

no, its not absolutely necessary; just convenient is all

OpenStudy (amistre64):

\[x = 7~sec(t)\] \[sec(t) = x/7\]